Circular No. 6280

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 Y1 S. Nakano, Sumoto, forwards the following additional positions:

1995 UT α_{2000} δ_{2000} Observer
Dec. 27.80634 14^{h}24^{m}56.^{s}02 −24^{o}32′02′′.9 Kushida
27.81328 14 24 57.50 −24 31 57.8  _{”}
27.81881 14 24 58.64 −24 31 53.8  _{”}
27.83973 14 25 03.44 −24 31 42.6 Kojima
27.84601 14 25 04.78 −24 31 39.0  _{”}
27.84832 14 25 05.33 −24 31 36.1  _{”}

Y. Kushida (Yatsugatake South Base Observatory). 0.40-m reflector +

CCD. Coma diameter ∼ 8′; strong condensation. T. Kojima (Chiyoda). 0.25-m reflector + CCD. Comet seems brighter than

on previous morning. Coma extension is now toward p.a. 250^{o}.

GLIESE 229B K. Matthews, T. Nakajima, S. Kulkarni, and B. Oppenheimer, Califor-

nia Institute of Technology, report: “We have obtained new measurements, at the Hale 5-m and the Keck 10-m telescopes, of Gliese 229B, the brown dwarf companion of Gliese 229A (reported by Nakajima et al. 1995, Nature 378, 463). Discrepancies in the K-band magnitude with the published value lead to a reanalysis of all the data. The published near-infrared photometric values (J, H, K_{s}, and K) are in error because of incorrect accounting for the scattered light of Gliese 229A. New measurements at longer wavelengths also have been made. For the aid of observers, the measurements of the flux (1σ uncertainties and upper limits) in a 2′′-diameter beam are as follows: Z [0.95–1.11 μm], magnitude 15.3, 1.5 mJy; J [1.15– 1.40 μm], 14.2, 3.2; H, [1.49–1.81 μm], 14.3, 2.0; K_{s}, [1.99–2.32 μm], 14.3, 1.3; K, [2.03–2.43 μm], 14.4, 1.2; [3.03–3.13 μm], 13.5, 1.5 ± 0.2; [3.28–3.34 μm], [14.4, <0.6; [3.38–3.45 μm], [14.7, <0.4; [3.67–3.97 μm], 11.8, 5.0 ± 1.0; [3.97–4.02 μm], 11.4, 6.6 ± 1.0; L′ [3.37–4.01 μm], 13.4, 1.1 ± 0.8; M [4.47–5.08 μm], 10.9, 7 ± 3; N [8.0–13.1 μm], 9.9, 4.2 ± 0.9. The uncertainties if not explicitly given are ± 10 percent. Oppenheimer et al. (1995, Science 270, 1478) have reported a low-resolution spectrum (range 1–2.5 μm) of Gliese 229B. The above magnitudes should be used to correct the normalization in each of the four near-infrared bands (Z, J, H, and K).”

1995 December 27 Daniel W. E. Green